### v2

@@
a School of Physical Sciences, and CAS Center for Excellence in Topological Quantum Computation, University of Chinese Academy of Sciences, Beijing 100049, China
@@ 1 Introduction
Weyl semimetals represent a novel topological system characterised by the presence of only a few band touching points within the Brillouin zone. These touching points are named as Weyl nodes, which always appear in pairs due to the no-go theorem[20, 21]. Accordingly, the fundamental Weyl semimetal is the system that contains only one pair of Weyl nodes, which is referred to as the ideal Weyl semimetal[22]. The two Weyl nodes exhibit distinct values of the Chern number of ±$\pm 1$, which represents different chirality. From an alternative perspective, the experiments demonstrate that the Weyl nodes can only be annihilated if two Weyl nodes with opposite chiralities merge together. This result demonstrates the robustness of the Weyl nodes and also confirms the findings of the calculation of the Chern number, which becomes zero when such two Weyl nodes meet together[18]. In light of this, the Chern number can be regarded as a topological invariant of a Weyl semimetal system, which can also be applied to systems that possess additional pairs of Weyl nodes, such as the $\boldmath{Z}_{2}$ Weyl semimetal, which features two pairs of Weyl nodes[23].
@@
AdS/CFT, which is based on the holographic principle of gravity, demonstrates the duality between the classical gravity and a strongly coupled field theory. It represents an effective methodology for the study of strongly coupled many-body systems[29, 30, 31]. A number of topological semimetals have been the subject of investigation in AdS/CFT, including Dirac semimetals[32], Weyl semimetals[5, 6], nodal line semimetals[33, 34] and $\boldmath{Z}_{2}$ Weyl semimetals[23]. Furthermore, the coexistence of two distinct types of topological semimetals has also been investigated[35]. These studies contribute to a more profound comprehension of topological systems and gravity.
@@
In this paper, we first review the procedure for the calculation of the topological invariant for the holographic Weyl semimetal and perform numerical calculations for the topological invariants in parameter regimes where backreactions of matter fields are important, which generalizes the previous probe limit calculation in [36] to more general circumstances. We examine the fermionic spectral functions of probe fermions on the holographic Weyl semimetal utilizing a numerical method and reveal the existence of multiple Weyl nodes. This is a special feature of holographic systems with backreactions of the matter fields, which does not show up in the probe limit Weyl semimetal calculated in [36]. We analyze the effective band structures obtained from the topological Hamiltonian and calculate the topological invariants for these nodes using numerics, which turn out to be ±$\pm 1$, indicating the topologically nontrival nature of the Weyl nodes in holographic Weyl semimetals.
@@
We then proceed to analyze the effective band structures and calculate the topological invariants for the $\mathbb{Z}_{2}$ Weyl semimetal utilizing the topological Hamiltonian method. The definition of the topological number of a $\mathbb{Z}_{2}$ Weyl semimetal differs from that of a typical Weyl semimetal, thus necessitating modification. This is because, in order to make clear the topology of a $\mathbb{Z}_{2}$ Weyl semimetal which possesses two pairs of nodes, it is not sufficient to consider solely the chirality of the nodes. Given that the degree of freedom of the spin has been introduced in $\mathbb{Z}_{2}$ Weyl semimetals, it is logical to conclude that the spin Chern number, a topological invariant that has been used to describe the spin Hall effect in numerous many-body systems[14], is the appropriate choice for characterising the topological properties of $\mathbb{Z}_{2}$ Weyl semimetals.
@@
In light of these considerations, we first undertake a review of the topological band theory, focusing on the physical and geometric significance of the Chern number and facilitating a more intuitive understanding of both the Chern number and the spin-Chern number. We subsequently revisit the effective field theory model[38] and the holographic model of the $\mathbb{Z}_{2}$ Weyl semimetal[23]. The corresponding topological invariant is calculated using the topological Hamiltonian method in the holographic model of the $\mathbb{Z}_{2}$ Weyl semimetal and compared with the case in the effective model of weak coupling. Furthermore, the distribution of the Weyl points intrinsic to the holographic $\mathbb{Z}_{2}$ Weyl semimetal is illustrated through the spectrum of probe fermions. The outcomes demonstrate that the distributions of the aforementioned Weyl points bear both resemblance and difference with those observed in the momentum space of the weak effective field theory model.
@@
The paper is organized as follows. In section 2, a review of the topological invariant for the Weyl semimetal is provided. New numerical calculations for topological invariants of holographic Weyl semimetal are presented. In section 3, the effective field theory model and the holographic model of the $\boldmath{Z}_{2}$ Weyl semimetal are reviewed, respectively. The effective band structure for the holographic $\boldmath{Z}_{2}$ Weyl semimetal is obtained through probe fermions, with the requisite zero-frequency Green’s functions being derived for the subsequent topological Hamiltonian calculations. In section 4, the topological Hamiltonian method was employed to calculate the corresponding topological invariant of the holographic $\boldmath{Z}_{2}$ Weyl semimetal. Section 5 is devoted to conclusions and open questions.
@@ 2.1 Basics on topological invariant for Weyl semimetals
Consider a weak coupling system which is described by a Hamiltonian $\hat{H}(k)$, at every $k$ in momentum space, the eigenvalue $E_{n}(k)$ and corresponding eigenvector |$\ket{n(k)}$ can be obtained by solving the eigen-function of $\hat{H}(k)$. Take only one branch of eigenvalue $E_{n}(k)$ and the corresponding eigenvector |$\ket{n(k)}$ as an example, we have
$\hat{H}(k)\ket{n(k)}=E_{n}(k)\ket{n(k)}.$ (2.1)
The corresponding Berry connection $\omega$ and Berry curvature $\Omega$ can be calculated using |$\ket{n(k)}$
$\omega=i\braket{n(k)}{d}{n(k)},~{}\Omega=id\braket{n(k)}{\wedge d}{n(k)}.$ (2.2)
@@
$c_{1}(E_{n})=-\frac{[\Omega]}{2\pi i}=-\frac{[d\braket{n(k)}{\wedge d}{n(k)}]}{2\pi}.$ (2.3)
Notice that $[\Omega]$ is the class which Berry curvature $\Omega$ represents in de Rham cohomology group, i.e. for an arbitrary exact 2-form $\eta$, $[\Omega]=[\Omega+\eta]$. When we need to integrate $[\Omega]$ over a closed surface, we just integrate Berry curvature $\Omega$ over the surface, for the integration of any exact form over a closed surface is 00. Now given a 2-dimensional close surface $M$ in momentum space, we can integrate the Chern class $c_{1}$ all over $M$ to get the Chern number $C$ corresponding to energy band $E_{n}$
$C(E_{n})=\int_{M}c_{1}=-\int_{M}\frac{d\braket{n(k)}{\wedge d}{n(k)}}{2\pi}.$ (2.4)
@@ 2.2 Topological invariant in effective field theoretical model for Weyl semimetal
The left panel of Fig. 1 illustrates the eigenenergy, otherwise termed the dispersion relation (2.10), of the Hamiltonian (2.9) in relation to $k_{z}$. Without loss of generality, it is assumed that $b_{z}=2$ and $m=1$ in (2.10). Two band crossing points are formed and located at $k_{z}=\pm\sqrt{3}$, respectively. It can be posited that the aforementioned two crossing points can be regarded as two Weyl nodes, which can in principle be obtained from a single Dirac point. The low-energy excitation in the vicinity of the Weyl node can be described using the Weyl equation, which implies that the elementary excitation is a Weyl fermion. Each of the nodes possesses a unique chirality, which can be distinguished by the topological number, as can be obtained by integrating the Berry curvature, (2.3), all over a closed surface. As illustrated in the right-hand panel of Fig. 1, the corresponding Berry curvature is demonstrated. It can be observed that the two Weyl nodes act as a sink and a source for the Berry curvature, respectively, thus indicating that these two nodes possess opposite values of the topological number. The result indicates that the topological number for the red point is $+1$, while the blue point possesses a negative topological number. The classification of two Weyl nodes is dependent upon their topological number.
@@
As a result, the topological invariant, can be designated as a Weyl charge, which bears resemblance to electric charge. The Weyl charge surrouneded by closed surface $M$ is can be defined as
@@ 2.3 Topological invariant in holographic Weyl semimetal: the topological Hamiltonian method
In order to circumvent the aforementioned limitations, an effective Hamiltonian need to be introduced for the system. This is to be defined by the zero frequency Green’s function, and it has been shown to comprise all the topology information of the system. This has been termed the “topological Hamiltonian”[26, 27, 28]. In a sense, the zero frequency Green’s function $G(0,k)$ contains all the information of $G(\omega,k)$ for $\omega$ is arbitrary value. All the poles of $G(\omega,k)$ locates on the $\omega=0$ plane, just like that a meromorphic function whose poles all locates at real axis. It is widely known that the pole contains all the information of the function, i.e. the near real axis part of the function can confirm the whole function. In this subsection, the definition of the topological Hamiltonian is provided. Subsequently, the ideal Weyl semimetal case is employed as a demonstrative exemplar to illustrate the calculation of the topological invariant in the holographic model by the topological Hamiltonian.
@@ 2.3.1 The topological Hamiltonian method
Although the topological Hamiltonian is not the exact Hamiltonian of the strong coupling system (there is not a Hamiltonian can be used to describe the whole strong coupling system), the topological Hamiltonian can be used to give the exact topological invariant, which cannot be calculated from the quasiparticle effective Hamiltonian. In conclusion, as long as ther Green’s function $G(\omega,k)$ does not have any poles besides on the $\omega=0$ plane, we can treat the topological Hamiltonian as a Hamiltonian in weak coupling system to calculate the topological invariant as it has been stated in previous works[41].
@@ 2.3.2 Review of the holographic Weyl semimetal
[…]
where $V$ is a vector $U(1)$ gauge field and $A$ is the axial $U(1)$ gauge field. $F_{V},F_{A}$ are the corresponding field strength, respectively. The coefficient $\alpha$ corresponds to the Chern-Simons term and is associated with the chiral anomaly. The scalar field $\Phi$ is charged only under the axial gauge symmetry. The corresponding covariant derivative is defined as $D_{a}:=\partial_{a}-iqA_{a}$. $R$ is the scalar curvature. The scalar bulk mass is set to be $m^{2}=-3$, and the radius of the AdS $L=1$ without loss of generality. We will choose other parameters in the (2.15) to be $2\kappa^{2}=1,\alpha=1,\lambda=\frac{1}{10}$ and $q=1$ without loss of generality.
@@ 2.3.3 Green’s functions from probe fermions
’ Then using the series definition of Bessel K function, we can investigate the behaviour of the solution near the boundary
@@ 2.3.4 Basic procedure for topological invariant of holographic Weyl semimetal
where $k_{l}:=(k_{x},k_{y},k_{z}+(-1)^{l}a_{1}),~{}l=1,2$, and $\cdots$ denotes higher order terms.
@@
The calculation of the holographic Green’s function and the consequent topological invariant for the holographic Weyl semimetal in the limit $M/b\to 0$(corresponding to the case where no back reaction of $A_{z}$ was considered) and for small $M/b$($\frac{M}{b}\ll\frac{M}{b}_{(c)}$, corresponding to a small back reaction) have been obtained in [36] using an analytic or semi-analytic method. However, for general $\frac{M}{b}$, where the back reaction becomes large, the analytical method becomes invalid. The present work employs numerical methods to facilitate the calculation of instances and the results of this calculation are presented in subsection 2.3.5. The calculation process is generally comprised of the following three steps.
The initial step in this process is to produce the numerical background for the holographic Weyl semimetal. The second step is to calculate the solutions of probe fermions on the holographic Weyl semimetal background numerically. We pick a very small $r_{0}>0$ and set the in-falling boundary condition for $\psi_{horizon}$. Under the chiral representation, the obtained numerical results of $\Psi_{1},\Psi_{2}$ are then utilised to determine the coefficients ${a_{1}}^{i}$ and ${a_{2}}^{i}$ in (2.38).
@@ 2.3.5 Fermion spectral functions and topological invariant for holographic Weyl semimetal
With infinite many nodes, it is important to use the topological invariant to detect which are trivial ones and which are nontrivial ones. Trivial topological invariant for a Weyl node indicates an accidental band crossing which is not topologically protected, i.e. it could become a gap under a small perturbation. Nontrivial topological invariant, which is ±$\pm 1$ for a Weyl node, denotes a band crossing that is topologically protected and cannot be gapped under small perturbations. For a pair of Weyl nodes, their topological invariants are opposite with each other. Therefore, it is also important to indicate which two Weyl nodes are a pair of nodes that could be annihilated to form a Dirac node. This could be analyzed by studying the band structures for a set of values of $\frac{M}{b}$ up to the Dirac critical point, which be presented in a future work[51].
Here we use the combined method of topological invariants and the observation of the band structure to distinguish Weyl nodes with opposite Weyl charges. The detailed analysis will be shown in the next section for the holographic $Z_{2}$ Weyl semimetal, which has the same structure as the Weyl nodes in the current holographic Weyl semimetal model. We summarize the result here. All the Weyl nodes in Figure. 2 have nontrivial topological charges of $+1$ or $-1$, consistent with the fact that all band crossings at the $k_{z}$ axis are not accidental touchings. Regarding which two Weyl nodes could be viewed as a pair of nodes with opposite charges, we could show from the result of the topological invariants that the two nearest adjacent ones have opposite Weyl charges. However, there is a subtle point here. As illustrated in Fig. 2, upon increasing the momentum $k_{z}$ from zero, the distances between a specific pair of Weyl nodes initially decrease to (almost)zero and subsequently increase rapidly. This phenomenon can be substantiated through the calculation of the results of the topology invariant, a subject that will be addressed in Sec. 4. As we could see from Figure. 2, this due to a band interchanging in the region between $k_{z}=0.6$ and $k_{z}=0.8$, and to the left of this region, the Weyl charges of the two nearest adjacent Weyl nodes display as $-1$,$+1$ while to the right of this region, the two charges are $+1$, $-1$ due to the band interchanging.

### v3

@@
a School of Physical Sciences, University of Chinese Academy of Sciences, Beijing 100049, China
@@ 1 Introduction
Weyl semimetals represent a novel topological system characterised by the presence of only a few band touching points within the Brillouin zone. These touching points are named as Weyl nodes, which always appear in pairs due to the no-go theorem[20, 21]. Accordingly, the fundamental Weyl semimetal is the system that contains only one pair of Weyl nodes, which is referred to as the ideal Weyl semimetal[22]. The two Weyl nodes exhibit distinct values of the Chern number of $\pm 1$, which represents different chirality. From an alternative perspective, the experiments demonstrate that the Weyl nodes can only be annihilated if two Weyl nodes with opposite chiralities merge together. This result demonstrates the robustness of the Weyl nodes and also confirms the findings of the calculation of the Chern number, which becomes zero when such two Weyl nodes meet together[18]. In light of this, the Chern number can be regarded as a topological invariant of a Weyl semimetal system, which can also be applied to systems that possess additional pairs of Weyl nodes, such as the Weyl-Z2 semimetal, which features two pairs of Weyl nodes[23].
@@
AdS/CFT, which is based on the holographic principle of gravity, demonstrates the duality between the classical gravity and a strongly coupled field theory. It represents an effective methodology for the study of strongly coupled many-body systems[29, 30, 31]. A number of topological semimetals have been the subject of investigation in AdS/CFT, including Dirac semimetals[32], Weyl semimetals[5, 6], nodal line semimetals[33, 34] and Weyl-Z2 semimetals[23]. Furthermore, the coexistence of two distinct types of topological semimetals has also been investigated[35]. These studies contribute to a more profound comprehension of topological systems and gravity.
@@
In this work, we study the topological invariants for a different system: the holographic Weyl-Z2 semimetal, which involves two sets of topological invariants: the Weyl charge and the Z2 charge. The calculation of the Weyl charge is similar to the calculation of the topological invariant for the pure holographic Weyl semimetal in [36]. However, in [36], to avoid heavy numerics only analytic calculations for the probe limit were performed, i.e. with no backreaction. Therefore, in this work, for the study of the more complicated holographic Weyl-Z2 semimetal system, we will first extend the calculation in [36] for the topological charge to holographic Weyl semimetal with backreaction using numerics. We will first review the procedure for the calculation of the topological invariant for the holographic Weyl semimetal and perform numerical calculations for the topological invariants in parameter regimes where backreaction of matter fields are important, which generalizes the previous probe limit calculation in [36] to more general circumstances. We examine the fermionic spectral functions of probe fermions on the holographic Weyl semimetal utilizing a numerical method and reveal the existence of multiple Weyl nodes. This is a special feature of holographic systems with backreaction of the matter fields, which does not show up in the probe limit Weyl semimetal calculated in [36]. We analyze the effective band structures obtained from the topological Hamiltonian and calculate the topological invariants for these nodes using numerics, which turn out to be $\pm 1$, indicating the topologically nontrival nature of the Weyl nodes in holographic Weyl semimetals.
@@
We then proceed to analyze the effective band structures and calculate the topological invariants for the Weyl-Z2 semimetal utilizing the topological Hamiltonian method. The definition of the topological number of a Weyl-Z2 semimetal differs from that of a typical Weyl semimetal, thus necessitating modification. This is because, in order to make clear the topology of a Weyl-Z2 semimetal which possesses two pairs of nodes, it is not sufficient to consider solely the chirality of the nodes. Given that the degree of freedom of the spin has been introduced in Weyl-Z2 semimetals, it is logical to conclude that the spin Chern number, a topological invariant that has been used to describe the spin Hall effect in numerous many-body systems[14], is the appropriate choice for characterising the topological properties of Weyl-Z2 semimetals.
@@
In light of these considerations, we first undertake a review of the topological band theory, focusing on the physical and geometric significance of the Chern number and facilitating a more intuitive understanding of both the Chern number and the spin-Chern number. We subsequently revisit the effective field theory model[38] and the holographic model of the Weyl-Z2 semimetal[23]. The corresponding topological invariant is calculated using the topological Hamiltonian method in the holographic model of the Weyl-Z2 semimetal and compared with the case in the effective model of weak coupling. Furthermore, the distribution of the Weyl points intrinsic to the holographic Weyl-Z2 semimetal is illustrated through the spectrum of probe fermions. The outcomes demonstrate that the distributions of the aforementioned Weyl points bear both resemblance and difference with those observed in the momentum space of the weak effective field theory model.
@@
The paper is organized as follows. In section 2, a review of the topological invariant for the Weyl semimetal is provided. New numerical calculations for topological invariants of holographic Weyl semimetal are presented. In section 3, the effective field theory model and the holographic model of the Weyl-Z2 semimetal are reviewed, respectively. The effective band structure for the holographic Weyl-Z2 semimetal is obtained through probe fermions, with the requisite zero-frequency Green’s functions being derived for the subsequent topological Hamiltonian calculations. In section 4, the topological Hamiltonian method was employed to calculate the corresponding topological invariant of the holographic Weyl-Z2 semimetal. Section 5 is devoted to conclusions and open questions.
@@ 2.1 Basics on topological invariant for Weyl semimetals
[…]
$\omega=i\braket{n(k)|d|n(k)},~\Omega=id\braket{n(k)|\wedge d|n(k)}.$ (2.2)
@@
$c_{1}(E_{n})=-\frac{[\Omega]}{2\pi i}=-\frac{[d\braket{n(k)|\wedge d|n(k)}]}{2\pi}.$ (2.3)
Notice that $\Omega]$ is the class which Berry curvature $\Omega$ represents in de Rham cohomology group, i.e. for an arbitrary exact 2-form $\eta$, $\Omega]=[\Omega+\eta]$. When we need to integrate $\Omega]$ over a closed surface, we just integrate Berry curvature $\Omega$ over the surface, for the integration of any exact form over a closed surface is 0. Now given a 2-dimensional closed surface $M$ in momentum space, we can integrate the Chern class $c_{1}$ all over $M$ to get the Chern number $C$ corresponding to energy band $E_{n}$
$C(E_{n})=\int_{M}c_{1}=-\int_{M}\frac{d\braket{n(k)|\wedge d|n(k)}}{2\pi}.$ (2.4)
@@ 2.2 Topological invariant in effective field theoretical model for Weyl semimetal
The left panel of Fig. 1 illustrates the eigenenergy, otherwise termed the dispersion relation (2.10), of the Hamiltonian (2.9) in relation to $k_{z}$. Without loss of generality, it is assumed that $b_{z}=2$ and $m=1$ in (2.10). Two band crossing points are formed and located at $k_{z}=\pm\sqrt{3}$, respectively. It can be posited that the aforementioned two crossing points can be regarded as two Weyl nodes, which can in principle be obtained from a single Dirac point. The low-energy excitation in the vicinity of the Weyl node can be described using the Weyl equation, which implies that the elementary excitation is a Weyl fermion. Each of the nodes possesses a unique chirality, which can be distinguished by the topological number, as can be obtained by integrating the Berry curvature, (2.3), all over a closed surface. As illustrated in the right-hand panel of Fig. 1, the corresponding Berry curvature is demonstrated. It can be observed that the two Weyl nodes act as a sink and a source for the Berry curvature, respectively, thus indicating that these two nodes possess opposite values of the topological number. The result indicates that the topological number for the red point is +1+1, while the blue point possesses a negative topological number. The classification of two Weyl nodes is dependent upon their topological number.
@@
As a result, the topological invariant, can be designated as a Weyl charge, which bears resemblance to electric charge. The Weyl charge surrouneded by closed surface $M$ can be defined as
@@ 2.3 Topological invariant in holographic Weyl semimetal: the topological Hamiltonian method
In order to circumvent the aforementioned limitations, an effective Hamiltonian need to be introduced for the system. This is to be defined by the zero frequency Green’s function, and it has been shown to comprise all the topology information of the system. This has been termed the “topological Hamiltonian”[26, 27, 28]. In a sense, the zero frequency Green’s function $G(0,k)$ contains all the information of $G(\omega,k)$ for $\omega$ takes an arbitrary value. All the poles of $G(\omega,k)$ are located on the $\omega=0$ plane, just like that a meromorphic function whose poles all locates at real axis. It is widely known that the pole contains all the information of the function, i.e. the near real axis part of the function can confirm the whole function. In this subsection, the definition of the topological Hamiltonian is provided. Subsequently, the ideal Weyl semimetal case is employed as a demonstrative exemplar to illustrate the calculation of the topological invariant in the holographic model by the topological Hamiltonian.
@@ 2.3.1 The topological Hamiltonian method
Although the topological Hamiltonian is not the exact Hamiltonian of the strong coupling system (there is not a Hamiltonian that can be used to describe the whole strong coupling system), the topological Hamiltonian can be used to give the exact topological invariant, which cannot be calculated from the quasiparticle effective Hamiltonian. In conclusion, as long as ther Green’s function $G(\omega,k)$ does not have any poles besides on the $\omega=0$ plane, we can treat the topological Hamiltonian as a Hamiltonian in weak coupling system to calculate the topological invariant as it has been stated in previous works[41]. [^1] Note that in the original proposal [26, 27, 28], for the topological Hamiltonian method to work, the regularity of the Green functions in the imaginary frequency space has to be required. Here because we have access to the more physical real frequency Green functions, we instead use the constraint that the real frequency Green functions are regular near the semimetal nodes, which are easy to be confirmed. Detailed analysis on this point is in Appendix D.
@@ 2.3.2 Review of the holographic Weyl semimetal
~~~- $(D_{a}\Phi)^{*}(D_{a}\Phi)-m|\Phi|^{2}-\frac{\lambda}{2}|\Phi|^{4}\bigg{]}\,$,
where $V$ is a vector $U(1)$ gauge field and $A$ is the axial $U(1)$ gauge field. $F_{V},F_{A}$ are the corresponding field strength, respectively. The coefficient $\alpha$ corresponds to the Chern-Simons term and is associated with the chiral anomaly. The scalar field $\Phi$ is charged only under the axial gauge symmetry. The corresponding covariant derivative is defined as $D_{a}:=\partial_{a}-iqA_{a}$. $R$ is the scalar curvature. The scalar bulk mass is set to be $m^{2}=-3$, and the radius of the AdS $L=1$ without loss of generality. We will choose other parameters in (2.15) to be $2\kappa^{2}=1,\alpha=1,\lambda=\frac{1}{10}$ and $q=1$ without loss of generality.
@@ 2.3.3 Green’s functions from probe fermions
Then using the series definition of Bessel K function, we can investigate the behaviour of the solution near the boundary
@@ 2.3.4 Basic procedure for topological invariant of holographic Weyl semimetal
where $k_{l}:=(-\omega,k_{x},k_{y},k_{z}+(-1)^{l}a_{1}),~l=1,2$, and $\cdots$ denotes higher order terms.
@@
The calculation of the holographic Green’s function and the consequent topological invariant for the holographic Weyl semimetal in the limit $M/b\to 0$, which corresponds to the case where no back reaction of $A_{z}$ was considered, and for small $M/b$, i.e. $\frac{M}{b}\ll\frac{M}{b}_{(c)}$, which corresponds to a small back reaction, have been obtained in [36] using an analytic or semi-analytic method. However, for general $\frac{M}{b}$, where the backreaction becomes large, the analytical method becomes invalid. If backreaction is not taken into account, the system would be in the $M/b\to 0$ limit, then several interesting features of the effective band structure would be lost:
  • There will only be one pair of Weyl nodes as in Fig.4, which is much similar to that in weak coupling system.
  • The two gapped bands would seem to have a crossing as in [36], which is an artifact of the $M/b\to 0$ limit.
  • Multiple Weyl nodes and the band inversion behavior - features that will be shown to exist in the backreaction case - would be lost.
Therefore, the present work employs numerical methods to facilitate the calculation of backreaction cases and the results of this calculation are presented in subsection 2.3.5. The calculation process is generally comprised of the following three steps.
The initial step in this process is to produce the numerical background for the holographic Weyl semimetal. The second step is to calculate the solutions of probe fermions on the holographic Weyl semimetal background numerically. We pick a very small $r_{0}>0$ and set the in-falling boundary condition for $\psi_{horizon}$. Before proceeding, it is necessary to clarify the physical meaning of $r_{0}$. The Lifshitz horizon in (2.36) is situated at the point of $r=0$. However, in order to solve the equations of motion numerically, it is necessary to employ an IR cutoff, designated as $r_{0}$, in order to import the initial condition of the Lifshitz IR geometry, i.e. the irrelevant deformations in the IR to flow the geometry to a required UV geometry. Then in accordance with the chiral representation, the obtained numerical results of $\Psi_{1},\Psi_{2}$ are utilised to determine the coefficients ${a_{1}}^{i}$ and ${a_{2}}^{i}$ in (2.38).
@@ 2.3.5 Fermion spectral functions and topological invariant for holographic Weyl semimetal
With infinite many nodes, it is important to use the topological invariant to detect which are trivial ones and which are nontrivial ones. Trivial topological invariant for a Weyl node indicates an accidental band crossing which is not topologically protected, i.e. it could become a gap under a small perturbation. Nontrivial topological invariant, which is $\pm 1$ for a Weyl node, denotes a band crossing that is topologically protected and cannot be gapped under small perturbations. For a pair of Weyl nodes, their topological invariants are opposite with each other. Therefore, it is also important to indicate which two Weyl nodes are a pair of nodes that could be annihilated to form a Dirac node. This could be analyzed by studying the band structures for a set of values of $\frac{M}{b}$ up to the Dirac critical point, which be presented in a future work[51].
Here we use the combined method of topological invariants and the observation of the band structure to distinguish Weyl nodes with opposite Weyl charges. The detailed analysis will be shown in the next section for the holographic $Z_{2}$ Weyl semimetal, which has the same structure as the Weyl nodes in the current holographic Weyl semimetal model. We summarize the result here. All the Weyl nodes in Figure. 2 have nontrivial topological charges of +1+1 or 1-1, consistent with the fact that all band crossings at the $k_{z}$ axis are not accidental touchings. Regarding which two Weyl nodes could be viewed as a pair of nodes with opposite charges, we could show from the result of the topological invariants that the two nearest adjacent ones have opposite Weyl charges. However, there is a subtle point here. As illustrated in Fig. 2, upon increasing the momentum $k_{z}$ from zero, the distances between a specific pair of Weyl nodes initially decrease to (almost)zero and subsequently increase rapidly. This phenomenon can be substantiated through the calculation of the results of the topology invariant, a subject that will be addressed in Sec. 4. As we could see from Figure. 2, this due to a band interchanging in the region between $k_{z}=0.6$ and $k_{z}=0.8$, and to the left of this region, the Weyl charges of the two nearest adjacent Weyl nodes display as 1-1,+1+1 while to the right of this region, the two charges are +1+1, 1-1 due to the band interchanging.